OpenStudy (anonymous):

Find an equation of a hyperbola with a=452 units and c=765 units. Assume that the transverse axis is horizontal. a.x^2/204,304 *y^2/380,921=1 b.x^2/380,921*y^2/204,304=1 c.x^2/452*y^2/313=1 d.x^2/313-y^2/452=1

OpenStudy (dumbcow):

well this is as simple as pythagorean thm a^2 +b^2 = c^2 --> b^2 = 765^2 - 452^2 hyperbola equation : \[\frac{x^{2}}{a^{2}}-\frac{y^{2}}{b^{2}} = 1\]